# Comparison Principles for Stochastic Volterra equations

### v1

@@ Appendix A Existence and uniqueness for Lipschitz coefficients
hold for a.a. $t\geq 0$ and all $x,y\in\mathbb{R}$ with some constant $C>0$. Then for each $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in[2,\infty]$, there exists a unique strong solution of the equation
(29) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}s.$
This solution satisfies $X\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L_{\mathrm{loc}}^{q}(\mathbb{R}_{+}))$ for each $p\in[2,q]$ with $p<\infty$. To be more precise, $\|X\|_{L^{p}(\Omega;L^{q}([0,T]))}\leq c$, where $c>0$ is a constant that only depends on $K,T,p,C$ [^3] The dependence on $C$ only comes through the growth estimate (28)..
@@
where $b_{\lambda}(s,x)=\mathrm{e}^{\lambda s}b(s,\mathrm{e}^{-\lambda s}x)$ and $\sigma_{\lambda}(s,x)=\mathrm{e}^{\lambda s}\sigma(s,\mathrm{e}^{-\lambda s}x)$. Conversely, let $Y_{\lambda}$ by a solution of (30), then $X^{\lambda}_{t}(g)=\mathrm{e}^{-\lambda t}Y_{\lambda}(t;g)$ satisfies (29). Therefore, it suffices to prove existence and uniqueness for (30). We solve (30) by a fixed point argument for the case $p=q$. Fix $T>0$, then $g|_{[0,T]}\in L^{q}([0,T]).$ For a given $\lambda<0$, define
@@
Suppose that $(g,K,b,\sigma)$ satisfies Assumption 2.2. Let $X$ be any solution of (29).
@@
    Under assumption 2.2.(i), $X-g$ has, for each $\theta<\gamma$, a continuous modification with locally $\theta$-Hölder continuous sample paths.
@@
    Under assumption 2.2.(ii), $X-g$ has, for each $\theta<\vartheta$ with
@@
Firstly, let us consider the case (ii). Since $q\geq\xi\left(1+\frac{2}{\eta}\right)$, we find $p\coloneqq\frac{\eta q}{(2+\eta)\xi}\geq 2$. Then, for $0<s<t\leq T,$
@@
where the finiteness of the expectation follows from Proposition A.1. Analogously we obtain from Itô ’s isometry, Jensen’s and Hölder’s inequality
@@
where $c>0$ is a constant that only depends on $p$,$\eta$, $C$, $T$ but not on $s$ or $t$, and the last equality follows from the particular form of $p$. Note that $\vartheta>0$ if and only if
@@
which is satisfied by Assumption 2.2. The Hölder continuity now follows from the Kolmogorov-Chentsov theorem. For the case of Assumption 2.2.(i), we let $p\geq 2$ be arbitrary. By following similar arguments to (32) and (33) and using $\operatorname*{ess\,sup}_{r\in[0,T]}\mathbb{E}[|X_{r}|^{p}]<\infty$ due to Proposition A.1, we find
$\mathbb{E}[|(X_{t}-g(t))-(X_{s}-g(s)|^{p}]\lesssim(t-s)^{p\gamma}.$
@@ Appendix B Splitting Method for Stochastic Volterra Equations
Given any solution $X$ of
$X_{t}=g(t)+\int_{0}^{t}K(t-s)[b(s,X_{s})\,\mathrm{d}s+\sigma(s,X_{s})\,\mathrm{d}B_{s}],$
where $g\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})\cap C((0,\infty))$ and $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ are measurable such that there exists $C\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ with
@@
the $\delta$-modulus of continuity of a function $f\colon[0,T]\longrightarrow\mathbb{R}$ and by $\|f\|_{[0,T]}\coloneqq\sup_{t\in[0,T]}|f(t)|$ its sup-norm. Below we extend the approximation [6, Lemma 3.1, Proposition 3.1] to the case of time-dependent coefficients and possibly singular $g$. As a first step, we prove a refined bound on the approximation where $g$ is regular.
@@
where we have used $\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s\leq\|C\|_{L^{2}([0,T]}^{2}$. By virtue of the discrete Gronwall inequality [13] and the standard inequality $1+x\leq\mathrm{e}^{x}$, we obtain $\sup_{1\leq k\leq N}\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}<\infty.$ Consequently, $\sup_{t\in[0,T]}\mathbb{E}[|\xi_{t}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}$ and so by construction of the approximation we arrive at (36). Let $t\in[t_{k-1},t_{k})$. We have $|\widehat{X}_{t}-\xi_{t}|^{2}\lesssim|\widehat{X}_{t}-\widehat{X}_{t_{k}-}|^{2}+|\widehat{X}_{t_{k}-}-\xi_{t}|^{2}$. For the second term, using (37) we arrive at
@@
By induction over $k$ we show that (41) holds on $[t_{k},t_{k+1})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k+1}\leq 0$. For $t\in(0,t_{1})$, (41) is evident and by assumption $\widetilde{x}_{1}=\frac{-f(t_{1})}{K(0_{+})}\leq 0$. Suppose (41) holds on $[t_{k-1},t_{k})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k}\leq 0.$ Then, we find for $t\in[t_{k},t_{k+1})$

### v2

@@ Appendix A Existence and uniqueness for Lipschitz coefficients
hold for a.a. $t\geq 0$ and all $x,y\in\mathbb{R}$ with some constant $C>0$. Then for each $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in[2,\infty]$, there exists a unique strong solution of the equation (1). This solution satisfies $X\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L_{\mathrm{loc}}^{q}(\mathbb{R}_{+}))$ for each $p\in[2,q]$ with $p<\infty$. To be more precise, $\|X\|_{L^{p}(\Omega;L^{q}([0,T]))}\leq c$, where $c>0$ is a constant that only depends on K,T,p,CK,T,p,C [^3] The dependence on $C$ only comes through the growth estimate (16)..
@@
where $b_{\lambda}(s,x)=\mathrm{e}^{\lambda s}b(s,\mathrm{e}^{-\lambda s}x)$ and $\sigma_{\lambda}(s,x)=\mathrm{e}^{\lambda s}\sigma(s,\mathrm{e}^{-\lambda s}x)$. Conversely, let $Y_{\lambda}$ by a solution of (17), then $X^{\lambda}_{t}(g)=\mathrm{e}^{-\lambda t}Y_{\lambda}(t;g)$ satisfies (1). Therefore, it suffices to prove existence and uniqueness for (17). We solve (17) by a fixed point argument for the case $p=q$.
Fix $T>0$, then $g|_{[0,T]}\in L^{q}([0,T]).$ For a given $\lambda<0$, define
@@
Suppose that $(g,K,b,\sigma)$ satisfies Assumption 2.1. Let $X$ be any solution of (1). Then the following assertions hold:
@@
    Under Assumption 2.1.(i), XgX-g has, for each $\theta<\gamma$, a continuous modification with locally $\theta$-Hölder continuous sample paths.
@@
    Under Assumption 2.1.(ii), XgX-g has, for each $\theta<\vartheta$ with
@@
Firstly, let us consider the case (ii). Since $q\geq 2\xi\left(1+\frac{2}{\eta}\right)$, we find $p\coloneqq\frac{\eta q}{(2+\eta)\xi}\geq 2$. Then, for $0<s<t\leq T,$
@@
where the finiteness of the expectation follows from Proposition A.1. Analogously, we obtain from Itô’s isometry, Jensen’s and Hölder’s inequality
@@
[…]
@@
which is satisfied by Assumption 2.1. The Hölder continuity now follows from the Kolmogorov-Chentsov theorem.
For the case of Assumption 2.1.(i), we let $p\geq 2$ be arbitrary. By following similar arguments to (19) and (20) and using $\operatorname*{ess\,sup}_{r\in[0,T]}\mathbb{E}[|X_{r}|^{p}]<\infty$ due to Proposition A.1, we find
$\mathbb{E}[|(X_{t}-g(t))-(X_{s}-g(s))|^{p}]\lesssim(t-s)^{p\gamma}.$
@@ Appendix B Splitting Method for Stochastic Volterra Equations
Given any solution $X$ of (1) where $g\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})\cap C((0,\infty))$ and $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ are measurable such that there exists $C\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ with
@@
the $\delta$-modulus of continuity of a function $f\colon[0,T]\longrightarrow\mathbb{R}$ and by $\|f\|_{[0,T]}\coloneqq\sup_{t\in[0,T]}|f(t)|$ its sup-norm. Below we extend the approximation [7, Lemma 3.1, Proposition 3.1] to the case of time-dependent coefficients and possibly singular $g$. As a first step, we prove a refined bound on the approximation where $g$ is regular.
@@
where we have used $\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s\leq\|C\|_{L^{2}([0,T]}^{2}$. By virtue of the discrete Gronwall inequality [13] and the standard inequality $1+x\leq\mathrm{e}^{x}$, we obtain $\sup_{1\leq k\leq N}\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}<\infty.$ Consequently, $\sup_{t\in[0,T]}\mathbb{E}[|\xi_{t}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}$ and so by construction of the approximation we arrive at (23).
Let $t\in[t_{k-1},t_{k})$. We have $|\widehat{X}_{t}-\xi_{t}|^{2}\lesssim|\widehat{X}_{t}-\widehat{X}_{t_{k}-}|^{2}+|\widehat{X}_{t_{k}-}-\xi_{t}|^{2}$. For the second term, using (24) we arrive at
@@
By induction over $k$ we show that (28) holds on $t_{k},t_{k+1})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k+1}\leq 0$. For $t\in(0,t_{1})$, (28) is evident and by assumption $\widetilde{x}_{1}=\frac{-f(t_{1})}{K(0_{+})}\leq 0$. Suppose (28) holds on $t_{k-1},t_{k})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k}\leq 0.$ Then, we find for $t\in[t_{k},t_{k+1})$